Find the vector equation of the line passing through 2i - 3j and i + j + k. Which point on this line has direction cosines, 1/√3, 1/√3, 1/√3 and why?
1
Expert's answer
2019-06-24T10:19:57-0400

Assume we have two vectors (a and b) with coordinates <2, -3, 0> and <1, 1, 1> in the (i, j, k) basis.

Firstly, we need to find a vector equation of the line passing through those points.



"t=w+\alpha v,"


where w is some vector with the end on that line and v is a vector parallel to it.

Let's assume

"w = a"

and

"v = a - b"

Then w has the coordinates <2, -3, 0> and v has <2 - 1, -3 - 1, 0 - 1> = <1, -4, -1>.

Vector equation of the line:


"t = <2,-3,0> + \alpha <1,-4,-1> = <2+\alpha ,-3-4\alpha,-\alpha>"

Next, find a vector equation of another line that has given direction cosines


"l = 1\/\u221a3, m = 1\/\u221a3, n = 1\/\u221a3"

Knowing the fact that


"l = x\/|r|, m = y\/|r|, n = z\/|r|,"

where |r| is a length of some vector on that line and x, y, z are its coordinates, we have


"x = |r|\/\u221a3, y = |r|\/\u221a3, z = |r|\/\u221a3"

and

"x = y = z"

So, vector equation of the line with given direction cosines is


"u = \theta + \beta <1, 1, 1> = <\beta , \beta , \beta >"

And the point we need is the intersection point of those lines


"<2+\alpha ,-3-4\alpha ,-\alpha >=<\beta ,\beta , \beta >"

"\beta = -\alpha , \alpha = -1, \beta = 1"

So, the point we need has the coordinates (1, 1, 1).

Answer. "t = <2,-3,0> + \alpha <1,-4,-1>; A = (1,1,1)" .